Find the axis of symmetry.
y = x^2 + x – 3

Answer by stanbon(75887)

Find the axis of symmetry.
y = x^2 + x – 3
----------------------
x^2+x+? = y+3+?
x^2+x+(1/2)^2 = y+3+(1/2)^2
(x+(1/2))^2 = y+(13/4)
-------------------
Vertex: (-1/2, -13/4)
Axis of symmetry: x = -1/2
====================
Cheers,
Stan H.